Can some one help me? Thank you

Solve the formula for the given letter. Assume all variables represent non negatibve numbers.
E=nc^2, for c

Answer by stanbon(75887)

Solve the formula for the given letter. Assume all variables represent non negative numbers.
E=nc^2, for c
--------------
c^2 = E/n
c = sqrt(E/n)
===============
Cheers,
Stan H.